# Dexterous Manipulation through Imitation Learning: A Survey

### v4

@@ Abstract
This work explores the intersection of IL and dexterous manipulation. With promising applications in manufacturing, healthcare, and home robotics, IL-based approaches allow robots to handle delicate objects, execute precise tasks, and operate in diverse, unstructured environments. However, key challenges remain—particularly in collecting high-quality demonstrations and enabling generalization from limited data, both of which are critical for realworld deployment. This survey aims to serve as a practical and accessible guide for practitioners seeking to apply IL-based methods to develop more capable and adaptable robotic systems in real-world scenarios.
@@ I Introduction
Over the past few decades, robotics has attracted intensive research interests, with dexterous manipulation emerging as a particularly popular focus. Dexterous manipulation aims to perform complex, precise, and flexible tasks (such as grasping an object, opening a drawer, and rotating a pencil) in various scenes with human-level dexterity using robotic hands or other end-effectors, as illustrated in Fig. 1. This high-precision manipulation capability supports a broad spectrum of applications, including industrial manufacturing [1, 2, 3, 4], space or underwater exploration, [5, 6, 7, 8] and medical care [9, 10, 11, 12]. Recently, the rapid development of imitation learning (IL) [13, 14], which seeks to acquire knowledge by observing and mimicking behaviors of humans or other agents, has led to notable advancements in computer graphics and robotics. As an intuitive approach to equip robots with human prior knowledge, especially in the ability to interact with objects and understand the scenes, IL has shown exceptional performance in enabling robots to perform tasks with human-like dexterity.
Figure 1: Examples of dexterous manipulation in the real world. Row 1: Dexterous manipulation platform used in our lab, Dextreme [15], DexCap [16], Row 2: Robotic telekinesis [17], Dexpilot [18], Row 3: Anyteleop [19].
Figure 2: Overview of dexterous manipulation based on imitation learning in this survey.
@@
However, when it comes to dexterous operations, such as assembling precision components or performing complex surgical procedures, a high degree of flexibility and multi-degree-of-freedom movement capabilities are required to execute complex, human-like tasks. The successful execution of dexterous manipulation hinges upon intricate and highly precise mechanical design, such as multi-fingered robotic hand [20] or anthropomorphic arms [21], as well as sophisticated control algorithms required to handle high-dimensional spaces [21] and multi-contact dynamics [22].
Recently, the exploration of employing IL in the field of robotics has garnered significant attention from researchers. Without the need for crafting complex world models and carefully designed reward functions, IL enables robots to learn tasks by observing and imitating expert demonstrations. This approach is intuitive, as the goal is for robots to substitute human labor by performing tasks like human experts. Specifically, the initial step involves collecting a dataset of expert demonstrations, which contains trajectories of manipulation tasks conducted by humans or well-trained agents. Robots use such trajectories as a reference to learn task behaviors. To ensure consistency, it is preferable to use identical robots during both the data collection and execution phases. However, this implementation does not facilitate data sharing among heterogeneous robotic systems. One solution is to map the trajectory of the original manipulator to the target robots, a process known as retargeting. Nonetheless, the process of humans operating robots for data collection remains time-consuming and labor-intensive in constructing large-scale datasets. To address this issue, researchers[23, 24, 17] have adopted pose estimation techniques from computer vision to develop mappings from human hands to robotic hands, effectively lowering the barriers of collecting demonstration data. Additionally, dataset augmentation enhances the ability to generalize to new objects and scenes, contributing to the expansion of the dataset.
@@
IL is particularly advantageous for dexterous manipulation tasks. This is because objects involved in dexterous manipulation are typically designed for human use, and thus, robots subject to these tasks are likely to have structures similar to humans or parts of the human body, such as humanoid robots, dual-arm manipulators, or dexterous hands. They usually require precise control, coordination, and adaptability, attributes that are challenging to achieve through traditional methods. Specifically, this learning paradigm includes various branches such as behavior cloning [25], [26], hybrid approaches (the combination of RL and IL) [27], [28], hierarchical IL [29], [30], and others, each contributing unique advantages to the learning process.
@@
Since the intersection of IL and dexterous manipulation represents a frontier in robotics research. In the past decade, several works including DAPG [27], which combines deep reinforcement learning(DRL) with human demonstrations to solve high-dimensional dexterous manipulation tasks; Implicit Behavioral Cloning [26] which focuses on improving robot policy learning from a mathematical perspective; Hiveformer [31], which explores creating multimodal interactive agents; Diffusion Policy [32] leverages recent advancement in generative models to achieve better performance in manipulation tasks, have been proposed and significantly expanded the boundaries of what is achievable in robotic dexterous manipulation. However, despite the recent considerable progress in this field, numerous challenges remain. Data collection for IL is labor-intensive and time-consuming [33], [16]. The acquisition of generalization ability from learned behaviors to new tasks and varying environments is also non-trivial [34]. Additionally, real-time control and sim-to-real transfer, where robots trained in the simulation must perform effectively in the real world, both hinder the application [35]. Addressing these challenges requires a concerted effort to develop more efficient data collection methods, improve learning algorithms, and enhance the physical capabilities of robotic systems.
@@
Building on significant advancements in the intersection of imitation learning (IL) and dexterous manipulation, several recent surveys have explored key aspects of this rapidly evolving field. Zare et al. [36] provide an extensive overview of the applications, progress, and challenges of imitation learning in both robotics and artificial intelligence. Arora and Doshi [37] focus on inverse reinforcement learning (IRL), particularly the process of inferring reward functions from observed behavior. Han et al. [38] delve into deep reinforcement learning (DRL), reviewing recent methods designed to optimize DRL algorithms for real-world applications. Li et al. [39] offer a comprehensive review of data collection methods and skill learning frameworks for robotic dexterous manipulation, emphasizing the key challenges in this area. Pitkevich and Makarov [40] address the Sim-to-Real transfer problem in DRL, discussing the deployment of robots trained in simulation to real-world environments. Welte and Rayyes [41] examine the potential of interactive imitation learning for humanoid robots, with a focus on how these methods can be transplanted for dexterous tasks. Finally, Tsuji et al. [42] provide a systematic review of imitation learning techniques for contact-rich tasks, discussing the unique challenges and emerging trends in this field. In contrast to these existing works, which primarily focus on specific aspects of imitation learning or dexterous manipulation, this survey aims to provide a comprehensive overview of IL-based dexterous manipulation approaches.
@@ II-A Dexterous Manipulation
In the field of robotics, dexterous manipulation [43, 44, 45] refers to the capability of robotic systems to execute intricate and precise tasks. These tasks often employ grippers or dexterous hands to grasp, maneuver, and manipulate objects [46]. Characterized by high degrees of freedom and fine motor skills, dexterous manipulation extends beyond simple pick-and-place operations to include activities such as tool use, object reorientation, and complex assembly tasks. Achieving such manipulation commonly involves using sophisticated end-effectors designed to emulate the versatility and finesse of human hands, such as multi-fingered hands or anthropomorphic robotic arms.
Dexterous manipulation poses several significant challenges, including precise control, high-dimensional motion planning, and real-time adaptability to dynamic environments [47]. The intricacies of these tasks demand robust mechanical design and advanced control algorithms capable of handling the complexities of multi-contact interactions and the variability inherent in real-world scenarios.
Traditional model-based methods [48, 49] have become inadequate for robots performing complex tasks due to the increasing complexity of manipulation tasks. Consequently, extensive research has been dedicated to learning-based approaches, with RL emerging as an effective method. Various works have exploited RL for robots to learn dexterous policies [50, 51, 52, 53, 54]. However, pure RL has several inherent drawbacks. RL algorithms often struggle with exploring high-dimensional action spaces efficiently in dexterous manipulation tasks [55]. Additionally, designing reasonable reward functions is challenging; flawed reward functions can affect exploration and learning speed, leading to inferior performance.
Recently, advancements in imitation learning (IL) have opened new avenues for addressing these challenges.
@@ II-B Imitation Learning
The main purpose of IL is to enable agents to learn and perform behaviors by imitating expert demonstrations [35]. In contrast, pure RL requires carefully designed reward functions and is particularly effective in scenarios where the desired behavior is difficult to describe in algorithms but can be easily demonstrated. IL employs these expert demonstrations to guide the learning process of agents by establishing the correlation between observed states and corresponding actions. Through IL, agents can transcend merely replicating basic and predefined behaviors within controlled and constrained environments, enabling them to autonomously execute optimal actions in complex, unstructured environments [56]. Consequently, IL significantly alleviates the burden on experts, facilitating efficient skill transfer.
Methodologies of IL can be broadly categorized into several sub-classes, including behavior cloning [57], inverse reinforcement learning (IRL) [58], and generative adversarial imitation learning (GAIL) [59]. Behavior cloning directly maps observed actions to the agent’s actions through SL techniques. Conversely, IRL aims to deduce the underlying reward structure that motivates the demonstrator’s behavior, allowing the agent to optimize its actions accordingly. GAIL employs adversarial training techniques to improve the imitation policy by distinguishing between expert and agent actions, thus refining the agent’s ability to replicate the desired behavior accurately.
@@ II-C Connections to Human Motor Learning
IL is fundamentally motivated by the way humans acquire motor skills through observation, demonstration, and practice. From a cognitive perspective, Bandura’s social learning theory [60] emphasizes that behavior can be learned by observing others, rather than solely through reinforcement or trial-and-error. This insight has inspired IL paradigms that rely on demonstration-based policy generation instead of environment-driven exploration.
At the neural level, the discovery of mirror neurons in primates provides biological evidence that action observation and execution may share overlapping representations [61]. These findings suggest that learning by watching others perform a task can activate internal motor plans, a mechanism analogous to behavioral cloning in robotics. Furthermore, models of human motor control, such as the internal model theory [62] and optimal feedback control [63], offer insights into how the brain predicts, plans, and corrects motor behaviors. These principles have influenced the design of IL algorithms that incorporate predictive models or closed-loop feedback refinement. For example, Nah et al. [64] present a modular robot control framework using Dynamic Movement Primitives (DMP) and Elementary Dynamic Actions (EDA), employing dynamical systems to predict and correct motor behaviors in a manner similar to internal models in the brain.
@@ II-D Long-horizon Dexterous Manipulation
Long-horizon dexterous manipulation involves executing complex tasks that require both fine-grained control and long-term planning. These tasks typically span long time horizons and require coherent, efficient execution of sequential actions. Without appropriate structure, such tasks can overwhelm agents, complicating both planning and execution. Task decomposition addresses this challenge by enabling agents to focus on different task components at various levels of abstraction, thereby facilitating the management of long-horizon manipulation [65]. In the context of IL, task decomposition is commonly realized through temporal abstraction, often implemented via techniques such as option discovery [66] or hierarchical policy structures [67, 68].
Temporal abstraction enables an agent to make decisions through temporally extended action sequences, thereby accelerating learning and simplifying complex tasks. Option discovery refers to the automated process of identifying and constructing these temporal abstractions, known as options, that comprise an internal low-level policy, an initiation set, and a termination condition. The options framework [69, 70] generalizes the traditional action concept by treating options as temporally extended actions, allowing planning over multiple time scales. The option-critic architecture [71] and its variants are widely adopted to automatically extract options from demonstration data, which are then scheduled by high-level policies to effectively organize the temporal structure of tasks. For example, DDCO [72] autonomously discovers and learns hierarchical options within continuous action spaces from expert demonstrations, enabling unsupervised task decomposition and hierarchical imitation learning for robotic control.
Hierarchical reinforcement learning is often integrated with IL to tackle long-horizon tasks and complex policy learning [73]. The fundamental principle involves leveraging human demonstrations to decompose tasks along temporal or functional dimensions, which reduces the policy search space and enhances both learning efficiency and generalization. For example, SRT-H [74] proposes a hierarchical language-conditioned IL framework that integrates high-level instruction parsing with low-level dexterous control to allow the execution of end-to-end autonomous surgical tasks.
@@ II-E Multi-agent or Collaborative Dexterous Manipulation
To enable collaborative dexterous manipulation, Kim et al. [74] introduced the SRT-H framework, which employs a hierarchical policy integrating high-level language instructions with low-level trajectory imitation. In a dual-arm da Vinci suturing task, the framework achieves precise bimanual coordination over long horizons, relying solely on human demonstrations without any reward supervision. Building on this line of research, BUDS [21] introduced a “stabilize-then-act” role decomposition with a visual keypoint stabilization mechanism, attaining 76.9% task success on complex bimanual dexterous tasks without reward signals or RL fine-tuning. Extending further, Bi-DexHands [75] framed bimanual dexterous manipulation as a comprehensive suite of over 20 cooperative sub-tasks, providing the first large-scale benchmark for evaluating the scalability of multi-agent RL in physically coordinated manipulation. More recently, BiDexHD [76] advanced collaborative dexterous manipulation by integrating multi-task human demonstration data with teacher–student policy learning. This framework enables robots to efficiently acquire a wide range of cooperative bimanual skills across numerous auto-constructed tasks, substantially improving task completion rates and zero-shot generalization.
@@
TABLE I: Comparison of different imitation learning approaches.
@@
Behavioral Cloning [77, 78, 20, 79, 14, 80, 81, 29, 77, 26, 82, 83, 84, 32, 85, 86] Supervised learning paradigm. Mapping from states to actions. No reward signals or exploration. Simple and easy to implement. Data-efficient with large
@@
[95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, Antotsiou2018TaskOrientedHM, 108]
@@
[109, 110, 111, 112, 113, 114, 115]
@@
[116, 117, 113, 118, 119, 120]

### v5

@@ Abstract
This work explores the intersection of IL and dexterous manipulation. With promising applications in manufacturing, healthcare, and home robotics, IL-based approaches allow robots to handle delicate objects, execute precise tasks, and operate in diverse, unstructured environments. However, key challenges remain—particularly in collecting high-quality demonstrations and enabling generalization from limited data, both of which are critical for real-world deployment. This survey aims to serve as a practical and accessible guide for practitioners seeking to apply IL-based methods to develop more capable and adaptable robotic systems in real-world scenarios.
@@ I Introduction
Over the past few decades, robotics has attracted intensive research interests, with dexterous manipulation emerging as a particularly popular focus. Dexterous manipulation aims to perform complex, precise, and flexible tasks (such as grasping an object, opening a drawer, and rotating a pencil) in various scenes with human-level dexterity using robotic hands or other end-effectors, as illustrated in Fig. 1. This high-precision manipulation capability supports a broad spectrum of applications, including industrial manufacturing [GONZALEZ2021283, 8629046, liang2017using, rebelo2014bilateral], space or underwater exploration, [diftler2012robonaut, brantner2021controlling, barbieri2018design, gharaybeh2019telerobotic], and medical care [zhang2020microsurgical, guo2019scaled, pugin2011history, talamini2002robotic]. Recently, the rapid development of imitation learning (IL) [peng2018deepmimic, zhang2018deep], which seeks to acquire knowledge by observing and mimicking behaviors of humans or other agents, has led to notable advancements in computer graphics and robotics. As an intuitive approach to equip robots with human prior knowledge, especially in the ability to interact with objects and understand scenes, IL has shown exceptional performance in enabling robots to perform tasks with human-like dexterity.
Figure 1: Examples of dexterous manipulation in the real world. Row 1: Customized dexterous manipulation platform, Dextreme [handa2023dextreme], DexCap [wang2024dex], Row 2: Robotic telekinesis [sivakumar2022robotic], Dexpilot [handa2020dexpilot], Row 3: Anyteleop [qin2023anyteleop].
Figure 2: Overview of imitation learning-based dexterous manipulation methods in this survey.
@@
However, when it comes to dexterous operations, such as assembling precision components or performing complex surgical procedures, a high degree of flexibility and multi-degree-of-freedom movement capabilities are required to execute complex, human-like tasks. The successful execution of dexterous manipulation hinges upon intricate and highly precise mechanical design, such as multi-fingered robotic hand [qin2022one] or anthropomorphic arms [grannen2023stabilize], as well as sophisticated control algorithms required to handle high-dimensional spaces [grannen2023stabilize] and multi-contact dynamics [zhu2023difflfd].
Recently, the exploration of employing IL in the field of robotics has garnered significant attention from researchers. Without the need to craft complex world models and carefully designed reward functions, IL enables robots to learn tasks by observing and imitating expert demonstrations. This approach is intuitive, as the goal is for robots to substitute human labor by performing tasks like human experts. Specifically, the initial step involves collecting a dataset of expert demonstrations, which contains trajectories of manipulation tasks conducted by humans or well-trained agents. Robots use such trajectories as a reference to learn task behaviors. To ensure consistency, it is preferable to use identical robots during both the data collection and execution phases. However, this implementation does not facilitate data sharing among heterogeneous robotic systems. One solution is to map the trajectory of the original manipulator to the target robots, a process known as retargeting. Nonetheless, the process of humans operating robots for data collection remains time-consuming and labor-intensive in constructing large-scale datasets. To address this issue, researchers [antotsiou2018task, li2019vision, sivakumar2022robotic] have adopted pose estimation techniques from computer vision to develop mappings from human hands to robotic hands, effectively lowering the barriers of collecting demonstration data. Additionally, dataset augmentation enhances the ability to generalize to new objects and scenes, contributing to the expansion of the dataset.
@@
IL is particularly advantageous for dexterous manipulation tasks. This is because objects involved in dexterous manipulation are typically designed for human use, and thus, robots subject to these tasks are likely to have structures similar to humans or parts of the human body, such as humanoid robots, dual-arm manipulators, or dexterous hands. They usually require precise control, coordination, and adaptability, attributes that are challenging to achieve through traditional methods. Specifically, this learning paradigm includes various branches such as behavior cloning [shi2023waypoint], [florence2022implicit], hybrid approaches (the combination of RL and IL) [rajeswaran2018learning], [ding2019goal], hierarchical IL [mandlekar2020learning], [belkhale2023hydra], and others, each contributing unique advantages to the learning process.
@@
Since the intersection of IL and dexterous manipulation represents a frontier in robotics research. In the past decade, several works including DAPG [rajeswaran2018learning], which combined deep reinforcement learning(DRL) with human demonstrations to solve high-dimensional dexterous manipulation tasks; Implicit Behavioral Cloning [florence2022implicit] which focused on improving robot policy learning from a mathematical perspective; Hiveformer [guhur2023instruction], which explored creating multimodal interactive agents; Diffusion Policy [chi2023diffusion] leveraged recent advancement in generative models to achieve better performance in manipulation tasks, have been proposed and significantly expanded the boundaries of what is achievable in robotic dexterous manipulation. However, despite the recent considerable progress in this field, numerous challenges remain. Data collection for IL is labor-intensive and time-consuming [zeng2021transporter], [wang2024dex]. The acquisition of generalization ability from learned behaviors to new tasks and varying environments is also non-trivial [haldar2023teach]. Additionally, real-time control and sim-to-real transfer, where robots trained in the simulation must perform effectively in the real world, both hinder the application [zare2023survey]. Addressing these challenges requires a concerted effort to develop more efficient data collection methods, improve learning algorithms, and enhance the physical capabilities of robotic systems.
@@
Building on significant advancements in the intersection of IL and dexterous manipulation, several recent surveys have explored key aspects of this rapidly evolving field. Zare et al. [zare2024survey] provided an extensive overview of the applications, progress, and challenges of imitation learning in both robotics and artificial intelligence. Arora and Doshi [ARORA2021103500] focused on inverse reinforcement learning (IRL), particularly the process of inferring reward functions from observed behavior. Han et al. [s23073762] delved into deep reinforcement learning (DRL), reviewing recent methods designed to optimize DRL algorithms for real-world applications. Li et al. [li2025developments] offered a comprehensive review of data collection methods and skill learning frameworks for robotic dexterous manipulation, emphasizing the key challenges in this area. Pitkevich and Makarov [pitkevich2024survey] addressed the Sim-to-Real transfer problem in DRL, discussing the deployment of robots trained in simulation to real-world environments. Welte and Rayyes [welte2025interactive] examined the potential of interactive imitation learning for humanoid robots, with a focus on how these methods can be transplanted for dexterous tasks. Finally, Tsuji et al. [tsuji2025survey] provided a systematic review of imitation learning techniques for contact-rich tasks, discussing the unique challenges and emerging trends in this field. In contrast to these existing works, which primarily focus on specific aspects of imitation learning or dexterous manipulation, this survey aims to provide a comprehensive overview of IL-based dexterous manipulation approaches.
@@ II-A Dexterous Manipulation
In the field of robotics, dexterous manipulation [salisbury1982articulated, Mason1985RobotHA, bai2014dex, liu2025fusion] refers to the capability of robotic systems to execute intricate and precise tasks. These tasks often employ grippers or dexterous hands to grasp, maneuver, and manipulate objects [okamura2000overview]. Characterized by high degrees of freedom and fine motor skills, dexterous manipulation extends beyond simple pick-and-place operations to include activities such as tool use, object reorientation, and complex assembly tasks. Achieving such manipulation commonly involves using sophisticated end-effectors designed to emulate the versatility and finesse of human hands, such as multi-fingered hands or anthropomorphic robotic arms.
Dexterous manipulation poses several significant challenges, including precise control, high-dimensional motion planning, and real-time adaptability to dynamic environments [yu2022dexterous]. The intricacies of these tasks demand robust mechanical design and advanced control algorithms capable of handling the complexities of multi-contact interactions and the variability inherent in real-world scenarios. Traditional model-based methods [buss1996dextrous, mordatch2012contact] have become inadequate for robots performing complex tasks due to the increasing complexity of manipulation tasks. Consequently, extensive research has been dedicated to learning-based approaches, with RL emerging as an effective method. Various works have exploited RL for robots to learn dexterous policies [popov2017data, williams1992simple, schulman2017proximal, dai2022analysing, xu2023dexterous, yin2023rot]. However, pure RL has several inherent drawbacks. RL algorithms often struggle with exploring high-dimensional action spaces efficiently in dexterous manipulation tasks [lillicrap2015continuous]. Additionally, designing reasonable reward functions is challenging; flawed reward functions can affect exploration and learning speed, leading to inferior performance. Recently, advancements in IL have opened new avenues for addressing these challenges.
@@ II-B Imitation Learning
The main purpose of IL is to enable agents to learn and perform behaviors by imitating expert demonstrations [zare2023survey]. In contrast, pure RL requires carefully designed reward functions and is particularly effective in scenarios where the desired behavior is difficult to describe in algorithms but can be easily demonstrated. IL employs these expert demonstrations to guide the learning process of agents by establishing the correlation between observed states and corresponding actions. Through IL, agents can transcend merely replicating basic and predefined behaviors within controlled and constrained environments, enabling them to autonomously execute optimal actions in complex, unstructured environments [ravichandar2020recent]. Consequently, IL significantly alleviates the burden on experts, facilitating efficient skill transfer.
Methodologies of IL can be broadly categorized into several sub-classes, including behavior cloning [bain1995framework], inverse reinforcement learning (IRL) [ng2000algorithms], and generative adversarial imitation learning (GAIL) [ho2016gener]. Behavior cloning directly maps observed actions to the agent’s actions through SL techniques. Conversely, IRL aims to deduce the underlying reward structure that motivates the demonstrator’s behavior, allowing the agent to optimize its actions accordingly. GAIL employs adversarial training techniques to improve the imitation policy by distinguishing between expert and agent actions, thus refining the agent’s ability to replicate the desired behavior accurately.
@@ II-C Connections to Human Motor Learning
IL is fundamentally motivated by the way humans acquire motor skills through observation, demonstration, and practice. From a cognitive perspective, Bandura’s social learning theory [bandura1977social] emphasized that behavior can be learned by observing others, rather than solely through reinforcement or trial-and-error. This insight has inspired IL paradigms that rely on demonstration-based policy generation instead of environment-driven exploration.
At the neural level, the discovery of mirror neurons in primates provides biological evidence that action observation and execution may share overlapping representations [rizzolatti2004mirror]. These findings suggest that learning by watching others perform a task can activate internal motor plans, a mechanism analogous to behavioral cloning in robotics. Furthermore, models of human motor control, such as the internal model theory [wolpert1995internal] and optimal feedback control [todorov2002optimal], offer insights into how the brain predicts, plans, and corrects motor behaviors. These principles have influenced the design of IL algorithms that incorporate predictive models or closed-loop feedback refinement. For example, Nah et al. [nah2025modular] presented a modular robot control framework using Dynamic Movement Primitives (DMP) and Elementary Dynamic Actions (EDA), employing dynamical systems to predict and correct motor behaviors in a manner similar to internal models in the brain.
@@ II-D Long-horizon Dexterous Manipulation
Long-horizon dexterous manipulation involves executing complex tasks that require both fine-grained control and long-term planning. These tasks typically span long time horizons and require coherent, efficient execution of sequential actions. Without an appropriate structure, such tasks can overwhelm agents, complicating both planning and execution. Task decomposition addresses this challenge by enabling agents to focus on different task components at various levels of abstraction, thereby facilitating the management of long-horizon manipulation [chen2025deco]. In the context of IL, task decomposition is commonly realized through temporal abstraction, often implemented via techniques such as option discovery [achiam2018variational] or hierarchical policy structures [sun2024hierarchical, wang2023hierarchical].
Temporal abstraction enables an agent to make decisions through temporally extended action sequences, thereby accelerating learning and simplifying complex tasks. Option discovery refers to the automated process of identifying and constructing these temporal abstractions, known as options, that comprise an internal low-level policy, an initiation set, and a termination condition. The options framework [sutton1999between, precup2000temporal] generalizes the traditional action concept by treating options as temporally extended actions, allowing planning over multiple time scales. The option-critic architecture [bacon2017option] and its variants are widely adopted to automatically extract options from demonstration data, which are then scheduled by high-level policies to effectively organize the temporal structure of tasks. For example, DDCO [krishnan2017ddco] autonomously discovers and learns hierarchical options within continuous action spaces from expert demonstrations, enabling unsupervised task decomposition and hierarchical IL for robotic control.
Hierarchical RL is often integrated with IL to tackle long-horizon tasks and complex policy learning [li2021hierarchical]. The fundamental principle involves leveraging human demonstrations to decompose tasks along temporal or functional dimensions, which reduces the policy search space and enhances both learning efficiency and generalization. For example, SRT-H [kim2025srt] proposed a hierarchical language-conditioned IL framework that integrates high-level instruction parsing with low-level dexterous control to allow the execution of end-to-end autonomous surgical tasks.
@@ II-E Multi-agent or Collaborative Dexterous Manipulation
To enable collaborative dexterous manipulation, Kim et al. [kim2025srt] introduced the SRT-H framework, which employs a hierarchical policy integrating high-level language instructions with low-level trajectory imitation. In a dual-arm da Vinci suturing task, the framework achieves precise bimanual coordination over long horizons, relying solely on human demonstrations without any reward supervision. Building on this line of research, BUDS [grannen2023stabilize] introduced a “stabilize-then-act” role decomposition with a visual keypoint stabilization mechanism, attaining 76.9% task success on complex bimanual dexterous tasks without reward signals or RL fine-tuning. Extending further, Bi-DexHands [chen2023bi] framed bimanual dexterous manipulation as a comprehensive suite of over 20 cooperative sub-tasks, providing the first large-scale benchmark for evaluating the scalability of multi-agent RL in physically coordinated manipulation. More recently, BiDexHD [zhou2024learning] advanced collaborative dexterous manipulation by integrating multi-task human demonstration data with teacher–student policy learning. This framework enables robots to efficiently acquire a wide range of cooperative bimanual skills across numerous auto-constructed tasks, substantially improving task completion rates and zero-shot generalization.
While task-level coordination addresses high-level planning, physical-level coordination must handle coupled, contact-rich dynamics inherent to dexterous manipulation. This challenge is addressed through unified model-based control. For example, Sleiman et al. [sleiman2021unified] proposed a whole-body MPC framework that jointly optimizes contact forces and centroidal dynamics in real time, enabling coherent coordination of locomotion and manipulation. Likewise, Yu et al. [yu2025generalizable] integrated global planning with adaptive MPC for dual-arm manipulation of deformable objects, capturing coupling through predictive dynamics. Beyond conventional hands, Zhong et al. [zhong2025paired] demonstrated that even micro-scale multi-agent coordination requires explicit handling of coupling dynamics, modeling nonlinear magnetic interactions between paired millirobots via data-driven inversion control and active disturbance rejection. These approaches collectively demonstrate that real-time compensation of coupled dynamics is essential for safe, responsive, and generalizable multi-agent IL policies.
@@
TABLE I: Comparison of different imitation learning approaches
@@
Behavioral Cloning [zhao2023learning, arunachalam2023dexterous, qin2022one, 8462901, zhang2018deep, osa2018algorithmic, ke2021grasping, mandlekar2020learning, zhao2023learning, florence2022implicit, shafiullah2022behavior, 9196935, chen2023diffusion, chi2023diffusion, ze2024d, ke20243d] Supervised learning paradigm. Mapping from states to actions. No reward signals or exploration. Simple and easy to implement. Data-efficient with large
@@
[tangkaratt2019vild, wang2021learning, zuo2021adversarial, antotsiou2021adversarial, liu2019hindsight, jiang2023mastering, wang2017robust, xiao2019wasserstein, arjovsky2017wasserstein, vahabpour2024diverse, tsurumine2022goal, shi2024ranking, zolna2021task, antotsiou2018task, tsurumine2019generative]
@@
[Kipf2018CompositionalIL, Xie2020DeepIL, 2409.16451, Xu2023XSkillCE, Wan2023LOTUSCI, wang2023mimicplay, 10610084]
@@
[Liang2024NeverEndingBA, Liu2023TAILTA, Wan2023LOTUSCI, Haldar2023PolyTaskLU, Gao2021CRILCR, Mete2024QueSTSS]